A point charge Q1 at x= 1 m and a charge Q2 at x= d. What is the relationship between these charges if the resulting electric field is zero at the following points: a) x= d/2 b) x= 2d c) x= -d/2?

(Answer: a) Q1/Q2= 1 b) Q1/Q2= -4 c) Q2/Q1= -9)


1
Expert's answer
2022-02-11T08:48:09-0500

a) Q1/Q2= 1

b) Q1/Q2= -4

c) Q2/Q1= -9